solve and check
x-12=-8(x-3)
------------
x^2-3x+5=x(x-8)+20
--------------
please help by 9 o'clock.
im begging you!!

Answer by rfer(16322)

x-12=-8(x-3)
x-12=-8x+24
9x=36
x=4
-----------------
x^2-3x+5=x(x-8)+20
x^2-3x+5=x^2-8x+20
5x=15
x=3